# Quasiparticle interference and spectral function of the UTe2 superconductive surface band

### v1

@@ IV.3 Surface spectral function
The surface spectral function $A_{s}(E,\mathbf{k}_{\parallel})$ is calculated from Eq. (22), and plotted at zero energy on Figs. 8 and 9 as a function of $k_{\#}$ and $k_{x}$ in the first Brillouin zone. For non-chiral pairings, we observe similar profiles of the surface spectral function for $A_{u}$, $B_{1u}$ and $B_{2u}$, with an amplitude which is of the same order compared to the one of the normal surface spectral function displayed on Fig. 3. On the contrary, for the $B_{3u}$ pairing, we obtain a strong enhancement in the amplitude of the surface spectral function compared to the normal one and compared to the other non-chiral pairings. This is the signature of the emergence of surface states at zero energy in the $B_{3u}$ case. Figure 9 shows that the amplitude is also enhanced for $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$, compared to the normal state, whereas it stays quite similar to the normal state for the other chiral pairings. Thus, both the density of states and the surface spectral function show the emergence of surface superconducting states for $B_{3u}$, $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$ pairings close to zero-energy.
@@
Figure 8: Surface spectral function as a function of $k_{\#}$ and $k_{x}$ at $E=0$ for non-chiral pairings. The gap value is fixed at $\Delta_{0}=0.3$ meV and the quasiparticle damping at $\eta=0.1$ meV. The strong intensity is coming from surface states indicating a topological character.
Figure 9: Same as in Fig. 8 for chiral pairings.
@@
To understand this in more detail, in Figs 10 and 11 we plot the surface band dispersion as a function of $k_{x}$ for $k_{\#}=-4.14$ nm-1, i.e. along one of the edges of the BZ. This clearly shows the emergence of sub-gap surface states centered at zero-energy for a $B_{3u}$ pairing, and very close to zero-energy for $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$ parings. All the other pairings exhibit subgap surface states centered at a finite energy of ±$\approx\pm 0.2$ meV. These results are summarized in Table 2. The full surface spectral function dispersion for the entire Brillouin zone is given in Figs. 26 and 27 in Appendix D.1, and is consistent with these observations. The observed surface states could potentially be topological surface states [46]. However, as this aspect is beyond the scope of the present work, we will refer to them simply as surface states.
@@
The choice for the damping value $\eta$ is constrained by the experiments [1]. By decreasing the value of $\eta$ in our simulation, the difference in amplitude between the superconducting and normal cases is accentuated, and the nodes presented schematically in Ref. [1] for $B_{2u}$ and $B_{3u}$ become visible in the surface spectral function (not shown here, but presented in the Methods section of Ref. [1]). We have checked that the surface states associated to these nodes do no contribute to the QPI scattering, and that the conclusions of the present analysis are unchanged by the value of $\eta$, and by the nodal point visibility.
@@
The symmetry constraints imposed by crystalline symmetries and time reversal on the edge contribution to the surface state are explained in details in Appendix B. The non-chiral pairings are mirror symmetric along both $k_{x}$ and $k_{\#}$ mirror axis, as in the bulk case, as shown in Fig. 8. These constraints arise from the $m_{x}$ and $\mathcal{T}m_{x}$ symmetries. For chiral pairings, we differentiate between those that preserve $m_{x}$ symmetry and those that do not, similar to the bulk case. The surface spectral function of the order parameter $A_{u}+iB_{3u}$ and $B_{1u}+iB_{2u}$ is mirror symmetric along the $k_{x}$ axis and protected by the $m_{x}$ symmetry. The surface spectral function for the other order parameters that break $m_{x}$ remains mirror symmetric along the $k_{\#}$ axis, a symmetry protected by $\mathcal{T}m_{x}$. These symmetry arguments hold for arbitrary energy $E$ (see Figs. 26 and 27 in Appendix D.1 for the plots at finite energy). At zero energy, the action of particle-hole symmetry leads to an enlarged symmetry group, making all superconducting order parameters symmetric along both the $k_{x}$ and $k_{\#}$ axis, as shown in Figs. 8 and 9.
@@ IV.4 Spin-resolved surface spectral function
The tight-binding model used in this work, described in Eq. (1), is spin $\uparrow$,$\downarrow$ , ↓ symmetric, with the two spin components completely decoupled. It means that any spin polarization of the spectral function comes from the triplet superconducting order parameter defined in Eq. (9). Figures 12 and 13 display the spin-resolved spectral function projected along the $\widehat{x}$-axis for both chiral and non-chiral pairings. In the non-chiral case, we plot this quantity only for $A_{u}$ and $B_{3u}$ pairings, as it cancels for $B_{1u}$ and $B_{2u}$ pairings. For the sake of simplicity we present the spin-resolved spectral functions along the $\widehat{y}$-axis and $\widehat{z}$-axis in Appendix C.1.
Since the normal state Hamiltonian is spin-independent, the bulk contribution results in a vanishing spin polarization for all pairings. On the contrary, the surface contribution can acquire a non-trivial spin-polarization depending on the form of the superconducting order parameter. From Figs. 1220 and 22, we see that the surface spectral function is polarized along $\widehat{x}$-axis for $B_{3u}$, along $\widehat{y}$-axis for $B_{2u}$ and along $\widehat{z}$-axis for $B_{1u}$, while the other pairings exhibit a more isotropic spin response (see Figs. 13,  21 and 23). In consequence, the $B_{3u}$ pairing case is the only one which has a spin-resolved surface spectral function with a polarization parallel to the $\widehat{x}$-axis, i.e. parallel to the (0-11) plane.
Figure 12: Spin-resolved surface spectral function along $\widehat{x}$-axis for $A_{u}$ and $B_{3u}$ pairings at $E=0$, $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV. Note that the spin-resolved surface spectral function along $\widehat{x}$-axis is equal to zero for $B_{1u}$ and $B_{2u}$ pairings.
Figure 13: Same as in Fig. 12 for chiral pairings.
@@ IV.5 QPI patterns
Figures 14 and 15 show the QPI patterns for the superconducting state of UTe2 described by various non-chiral and chiral OPs at zero energy. Once more we consider an impurity localized on the (0-11)-plane with impurity potential $U_{0}=0.2$ eV. There are two key differences compared to the QPI pattern obtained for UTe2 in the normal state (see Fig. 3): (i) the presence of a finite intensity along a horizontal band close to $q_{x}=0$ for all the pairings, which was absent in the normal state, and (ii) the presence at $q_{\#}=0$ and finite $q_{x}$ of an additional peak, denoted ${\bf q}_{1}$ for some of the OPs, especially $B_{3u}$, $A_{u}+iB_{1u}$, $A_{u}+iB_{2u}$, $A_{u}+iB_{3u}$, $B_{1u}+iB_{3u}$ and $B_{2u}+iB_{3u}$.
Figure 14: QPI patterns as a function of $q_{\#}$ and $q_{x}$ for non-chiral pairings at $E=0$, $U_{0}=0.2$ eV, $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV.
Figure 15: Same as in Fig. 14 for chiral pairings.
The various peaks in the QPI patterns are identified by the vectors ${\bf q}_{1}$ to ${\bf q}_{6}$. To help understanding the presence of these peaks, we have marked the scattering vectors associated to each of these peaks in Fig. 16, for both normal UTe2 and superconducting UTe2 with pairings $B_{2u}$ and $B_{3u}$. In the normal state case there are only two peaks, characterized by the scattering vectors ${\bf q}_{2}$ and ${\bf q}_{6}$ that are related to the crystal structure. For the $B_{2u}$ pairing two additional peaks are visible, characterized by vectors ${\bf q}_{4}$ and ${\bf q}_{5}$. Finally, for the $B_{3u}$ pairing, two other peaks are visible, characterized by vectors ${\bf q}_{1}$ and ${\bf q}_{3}$. Thus, at $E=0$, there are two peaks in the normal state, four peaks for $B_{2u}$, and six peaks for $B_{3u}$. The other non-chiral and chiral pairings show similar structure to either $B_{2u}$ and $B_{3u}$. We have identified the correspondence between the scattering vectors and the regions of high intensity in the surface spectral function. We note however that while some peaks would be expected from the spectral function analysis, their absence is related to destructive interference, leading to forbidden scattering processes between some of the bands.
It should be emphasized also that the QPI patterns depend slightly on the value of the impurity amplitude $U_{0}$ (see Eqs. (24) to (27)). However, we have checked that the main features and the conclusion of our analysis do not change when $U_{0}$ varies.
@@
The evolution of the QPI pattern with increasing energy $E$ from zero up to the value 0.25 meV is given in Figs. 28 and 29 in Appendix D.2. These figures indicate that the intensity of the QPI patterns reflects the intensity of the spectral function, and this is especially true for the $q_{1}$ peak whose maximum of intensity as a function of energy follows the maximum intensity in Figs 10 and 11 for each OP, as noted in Table 2.
We compare these results with the experimental observations in Ref. 1 presented in Figs. 17 and 18. In particular, the experimental results show a peak in the DOS at zero energy, as well as a non-dispersing ${\bf q}_{1}$ feature. Among the non-chiral states, that are most likely to describe the physics of UTe2, the $B_{3u}$ OP is the only one that has a ${\bf q}_{1}$ peak that is stable in position and remains visible up to $E\approx 0.15$ meV. This observation, combined with the fact that $B_{3u}$ is the only OP among the non-chiral one that exhibits a peak at zero energy in the DOS, indicates that $B_{3u}$ is the most likely candidate to describe the superconducting order parameter of UTe2.
Figure 17: Experimental results for (left panel) UTe2 DOS in the normal state (4.2 K) and in the superconducting state (0.28 K), and (right panel) QPI pattern at zero energy[1].
Figure 18: Experimental results for QPI pattern plotted in the first Brillouin zone, at increasing energy $E\in\{0,50,100,150,200,250\}$ $\mu$eV[1].
@@ V Discussions and conclusion
As outlined in the introduction, determining the pairing symmetry of UTe2 is challenging and requires a multi-method approach. Experimental data available concerning time-reversal symmetry breaking seems to indicate that a non-chiral pairing is the most probable hypothesis [13, 22]. Currently, there is active debate on whether the pairing symmetry is $B_{2u}$ or $B_{3u}$. Our analysis can help to resolve this debate by providing clear predictions to distinguish especially between the non-chiral pairing symmetries $A_{u}$, $B_{1u}$, $B_{2u}$, and $B_{3u}$ through STM experiments. Indeed, by comparing our results to the experimental ones[1], we can give strong arguments that the pairing state of superconducting UTe2 is $B_{3u}$. This is based notably on the existence of the ${\bf q}_{1}$ peak in the QPI patterns and on the presence of the zero-energy peak in the surface DOS.
If one takes into account also the chiral OPs, the $A_{u}+iB_{1u}$, $A_{u}+iB_{2u}$, $A_{u}+iB_{3u}$, $B_{1u}+iB_{3u}$ or $B_{2u}+iB_{3u}$ also show a ${\bf q}_{1}$ feature. However the three latter ones can be eliminated from the possible scenarios, since they have minimum in its DOS at zero-energy which contradicts the experimental results[1] as shown in the left panel of Fig. 17. The discrimination between the remaining pairings, $B_{3u}$, $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$, could be done through various other methods. As noted earlier, the chiral order parameters appear highly unlikely, a conclusion reinforced by the splitting of the zero-energy Andreev peak [21] under proximity with a $s$-wave superconductor, which suggests a time-reversal symmetric superconducting state. In the future spin resolved experiments may also be used to distinguish between these pairings, since we have shown in Section IV.4 that they have different spin-resolved spectral functions.
After our work was completed, we became aware of Ref. [50] which reports on a model for the band structure and the superconductive topological surface states of UTe2 using a different technique.

### v2

@@ IV.3 Surface spectral function
The surface spectral function $A_{s}(E,\mathbf{k}_{\parallel})$ is calculated from Eq. (22), and plotted at zero energy in Figs. 8 and 9 as a function of $k_{\#}$ and $k_{x}$ in the first Brillouin zone. For non-chiral pairings, we observe similar profiles of the surface spectral function for $A_{u}$, $B_{1u}$ and $B_{2u}$, with an amplitude which is of the same order compared with that of the normal surface spectral function displayed in Fig. 3. On the contrary, for the $B_{3u}$ pairing, we obtain a strong enhancement in the amplitude of the surface spectral function compared with the normal one and compared with the other non-chiral pairings. This is the signature of the emergence of surface states at zero energy in the $B_{3u}$ case. Figure 9 shows that the amplitude is also enhanced for $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$, compared with the normal state, whereas it stays quite similar to the normal state for the other chiral pairings. Thus, both the density of states and the surface spectral function show the emergence of surface superconducting states for $B_{3u}$, $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$ pairings close to zero-energy. In general, topological surface states form as arcs between the projection of the bulk nodes to the surface. Among the non-chiral OPs for which we have investigated the nodal structure, this seems to be the case only for $B_{3u}$. For the other pairings, the surface states are not connected to the nodal structure, and they form at finite energy; their origin is thus unclear.
@@
Figure 8: Surface spectral function as a function of $k_{\#}$ and $k_{x}$ at $E=0$ for non-chiral pairings. The gap value is fixed at $\Delta_{0}=0.3$ meV and the quasiparticle damping at $\eta=0.1$ meV.
Figure 9: Surface spectral function as a function of $k_{\#}$ and $k_{x}$ at $E=0$ for chiral pairings. The gap value is fixed at $\Delta_{0}=0.3$ meV and the quasiparticle damping at $\eta=0.1$ meV.
@@
To understand this in more detail, in Figs 10 and 11 we plot the surface band dispersion as a function of $k_{x}$ for $k_{\#}=-4.14$ nm-1, i.e. along one of the edges of the BZ, and along a close path ABCDA of the BZ in Appendix D.1. This clearly shows the emergence of sub-gap surface states centered at zero-energy for a $B_{3u}$ pairing, and very close to zero-energy for $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$ parings. All other pairings exhibit subgap surface states centered at a finite energy of $\approx\pm 0.2$ meV. These results are summarized in Table 2. The complete surface spectral function dispersion for the entire Brillouin zone is given in Figs. 31 and 32 in Appendix D.2, and is consistent with these observations. The observed surface states could potentially be topological surface states [sato2017topological]. However, as this aspect is beyond the scope of the present work, we refer to them simply as surface states.
@@
The choice of damping value $\eta$ is constrained by experiments [experimentalpaper]. By decreasing the value of $\eta$ in our simulation, the difference in amplitude between the superconducting and normal cases is accentuated, and the nodes presented schematically in Ref. [experimentalpaper] for $B_{2u}$ and $B_{3u}$ become visible in the surface spectral function (not shown here, but presented in the Methods section of Ref. [experimentalpaper]). We have checked that the surface states associated with these nodes do not contribute to the QPI scattering, and that the conclusions of the present analysis are unchanged by the value of $\eta$, and by the nodal point visibility.
@@
The symmetry constraints imposed by the crystalline symmetries and the time reversal on the edge contribution to the surface state are explained in detail in Appendix B. The non-chiral pairings are mirror symmetric along both $k_{x}$ and $k_{\#}$ mirror axis, as in the bulk case, as shown in Fig. 8. These constraints arise from the $m_{x}$ and $\mathcal{T}m_{x}$ symmetries. For chiral pairings, we differentiate between those that preserve $m_{x}$ symmetry and those that do not, similar to the bulk case. The surface spectral function of the order parameter $A_{u}+iB_{3u}$ and $B_{1u}+iB_{2u}$ is mirror-symmetric along the $k_{x}$ axis and protected by the $m_{x}$ symmetry. The surface spectral function for the other order parameters that break $m_{x}$ remains mirror symmetric along the $k_{\#}$ axis, a symmetry protected by $\mathcal{T}m_{x}$. These symmetry arguments hold for an arbitrary energy $E$ (see Figs. 31 and 32 in Appendix D.2 for the plots at finite energy). At zero energy, the action of particle-hole symmetry leads to an enlarged symmetry group, making all superconducting order parameters symmetric along both the $k_{x}$ and $k_{\#}$ axis, as shown in Figs. 8 and 9.
@@ IV.4 Spin-resolved surface spectral function
The tight-binding model used in this work, described in Eq. (1), is spin $\uparrow,\downarrow$ symmetric, with the two spin components completely decoupled. This means that any spin polarization of the spectral function comes from the triplet superconducting order parameter defined in Eq. (9). Figures 12 and 13 display the spin-resolved spectral function projected along the $\widehat{x}-$axis for both chiral and non-chiral pairings. In the non-chiral case, we plot this quantity only for $A_{u}$ and $B_{3u}$ pairings because it cancels for $B_{1u}$ and $B_{2u}$ pairings. For the sake of simplicity, we present the spin-resolved spectral functions along the $\widehat{y}-$axis and $\widehat{z}-$axis in Appendix C.1.
Since the normal state Hamiltonian is spin-independent, the bulk contribution results in a vanishing spin polarization for all pairings. On the contrary, the surface contribution can acquire a non-trivial spin-polarization depending on the form of the superconducting order parameter. From Figs. 1222 and 24, we see that the surface spectral function is polarized along $\widehat{x}-$axis for $B_{3u}$, along $\widehat{y}-$axis for $B_{2u}$ and along $\widehat{z}-$axis for $B_{1u}$, while the other pairings exhibit a more isotropic spin response (see Figs. 13,  23 and 25). In consequence, the $B_{3u}$ pairing case is the only one which has a spin-resolved surface spectral function with a polarization parallel to the $\widehat{x}-$axis, i.e. parallel to the (0-11) plane.
Figure 12: Spin-resolved surface spectral function along $\widehat{x}-$axis for $A_{u}$ and $B_{3u}$ pairings at $E=0$, $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV. Note that the spin-resolved surface spectral function along $\widehat{x}-$axis is equal to zero for $B_{1u}$ and $B_{2u}$ pairings. The symmetry of the spin texture can be understood from the structure of the superconducting order parameter for each pairing, as detailed in Appendix B.6.
Figure 13: Spin-resolved surface spectral function along $\widehat{x}-$axis for chiral pairings at $E=0$, $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV. Since chiral pairings have superconducting order parameter components along all spin directions, each chiral pairing exhibits non-zero spin polarization in every spin channel.
@@ IV.5 QPI patterns
Figures 14 and 15 show the QPI patterns for the superconducting state of UTe2 described by various non-chiral and chiral OPs at zero energy. Once more we consider an impurity localized on the (0-11)-plane with impurity potential $U_{0}=0.2$ eV. There are two key differences compared with the QPI pattern obtained for UTe2 in the normal state (see Fig. 3): (i) the presence of a finite intensity along a horizontal band close to $q_{x}=0$ for all the pairings, which was absent in the normal state, and (ii) the presence at $q_{\#}=0$ and finite $q_{x}$ of an additional peak, denoted ${\bf q}_{1}$ for some of the OPs, especially $B_{3u}$, $A_{u}+iB_{1u}$, $A_{u}+iB_{2u}$, $A_{u}+iB_{3u}$, $B_{1u}+iB_{3u}$ and $B_{2u}+iB_{3u}$.
Figure 14: QPI patterns as a function of $q_{\#}$ and $q_{x}$ for non-chiral pairings at $E=0$, $U_{0}=0.2$ eV, $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV. Because of their similar surface spectral functions, the $A_{u}$, $B_{1u}$ and $B_{2u}$ pairings exhibit relatively similar QPI patterns. In contrast, the $B_{3u}$ pairing displays an additional peak from from surface states scattering. The distinct nodal structures of the order parameters are not reflected in the QPI patterns due to the strong quasiparticle damping $\eta$.
Figure 15: QPI patterns as a function of $q_{\#}$ and $q_{x}$ for chiral pairings at $E=0$, $U_{0}=0.2$ eV, $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV. Similar to $B_{3u}$ pairing, $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$ display an additional peak from from surface states scattering. This peak is present but with weaker intensity for $A_{u}+iB_{3u}$. The distinct nodal structures of the order parameters are not reflected in the QPI patterns due to the strong quasiparticle damping $\eta$.
The various peaks in the QPI patterns are identified by the vectors ${\bf q}_{1}$ to ${\bf q}_{6}$. To help understand the presence of these peaks, we have marked the scattering vectors associated to each of these peaks in Fig. 16, for both normal UTe2 and superconducting UTe2 with pairings $B_{2u}$ and $B_{3u}$. In the normal state case, there are only two peaks, characterized by the scattering vectors ${\bf q}_{2}$ and ${\bf q}_{6}$ that are related to the crystal structure. For the $B_{2u}$ pairing two additional peaks are visible, characterized by vectors ${\bf q}_{4}$ and ${\bf q}_{5}$. Finally, for the $B_{3u}$ pairing, two other peaks are visible, characterized by vectors ${\bf q}_{1}$ and ${\bf q}_{3}$. Thus, at $E=0$, there are two peaks in the normal state, four peaks for $B_{2u}$, and six peaks for $B_{3u}$. The other non-chiral and chiral pairings show similar structure to either $B_{2u}$ and $B_{3u}$. We have identified the correspondence between the scattering vectors and the regions of high intensity in the surface spectral function. We note, however, that while some peaks would be expected from the spectral function analysis, their absence is related to destructive interference, leading to forbidden scattering processes between some of the bands.
It should also be emphasized that the QPI patterns depend slightly on the value of the impurity amplitude $U_{0}$ (see Eqs. (24) to (27)). However, we have checked that the main features and the conclusion of our analysis do not change when $U_{0}$ varies.
@@
The evolution of the QPI pattern with increasing energy $E$ from zero up to the value 0.25 meV is given in Figs. 33 and 34 in Appendix D.3. These figures indicate that the intensity of the QPI patterns reflects the intensity of the spectral function, and this is especially true for the $q_{1}$ peak whose maximum of intensity as a function of energy follows the maximum intensity in Figs 10 and 11 for each OP, as noted in Table 2.
We compare these results with the experimental observations in Ref. experimentalpaper presented in Figs. 17 and 18. In particular, the experimental results show a peak in the DOS at zero energy, as well as a non-dispersing ${\bf q}_{1}$ feature. Among the non-chiral states, that are most likely to describe the physics of UTe2, the $B_{3u}$ OP is the only one that has a ${\bf q}_{1}$ peak that is stable in position and remains visible up to $E\approx 0.15$ meV. This observation, combined with the fact that $B_{3u}$ is the only OP among the non-chiral one that exhibits a peak at zero energy in the DOS, indicates that $B_{3u}$ is the most likely candidate to describe the superconducting order parameter of UTe2.
Figure 17: Experimental measurement of differential conductance for (left panel) UTe2 in the normal state (4.2 K) and Andreev differential conductance in the superconducting state (0.28 K), and (right panel) QPI pattern at zero energy[experimentalpaper].
[…]
@@ V Discussions and conclusion
As outlined in the introduction, determining the pairing symmetry of UTe2 is challenging and requires a multi-method approach. Experimental data available concerning time-reversal symmetry breaking seems to indicate that a non-chiral pairing is the most probable hypothesis [gu2023detection, theuss2024single]. Currently, there is active debate on whether the pairing symmetry is $B_{2u}$ or $B_{3u}$. Our analysis can help to resolve this debate by providing clear predictions to distinguish especially between the non-chiral pairing symmetries $A_{u}$, $B_{1u}$, $B_{2u}$, and $B_{3u}$ through STM experiments. Indeed, by comparing our results to the experimental ones[experimentalpaper], we can give strong arguments that the pairing state of superconducting UTe2 is $B_{3u}$. This is based notably on the existence of the ${\bf q}_{1}$ peak in the QPI patterns and on the presence of the zero-energy peak in the surface DOS.
If one also takes into account the chiral OPs, the $A_{u}+iB_{1u}$, $A_{u}+iB_{2u}$, $A_{u}+iB_{3u}$, $B_{1u}+iB_{3u}$ or $B_{2u}+iB_{3u}$ also show a ${\bf q}_{1}$ feature. However, the latter three can be eliminated from the possible scenarios, since they have minimum in its DOS at zero-energy which contradicts the experimental results[experimentalpaper] as shown in the left panel of Fig. 17. The discrimination between the remaining pairings, $B_{3u}$, $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$, could be done using various other methods. As noted earlier, the chiral order parameters appear highly unlikely, a conclusion reinforced by the splitting of the zero-energy Andreev peak [gu2025pair] under proximity with a $s$-wave superconductor, which suggests a time-reversal symmetric superconducting state. In the future spin-resolved experiments may also be used to distinguish between these pairings, since we have shown in Section IV.4 that they have different spin-resolved spectral functions.
Note added. After our work was completed, we became aware of Ref. [christiansen2025nodal] which reports a model for the band structure and the superconductive topological surface states of UTe2 using a different technique.